- Press Release
- Sep 26, 2022
NASA Hubble Space Telescope Daily Report #4921
HUBBLE SPACE TELESCOPE DAILY REPORT #4921
Continuing to Collect World Class Science
PERIOD COVERED: 5am August 31 – 5am September 1, 2009 (DOY 243/09:00z-244/09:00z)
ACS CCD Monitoring and Calibration for WFC3
This program is a smaller version of our routine CCD monitoring program, designed to run throughout SMOV, after which our regular Cycle 17 CAL proposal will begin. This program obtains the bias and dark frames needed to generate reference files for calibrating science data, and allows us to monitor detector noise and the growth of hot pixels.
CCD Daily Monitor (Part 1)
This program comprises basic tests for measuring the read noise and dark current of the ACS WFC and for tracking the growth of hot pixels. The recorded frames are used to create bias and dark reference images for science data reduction and calibration. This program will be executed four days per week (Mon, Wed, Fri, Sun) for the duration of Cycle 17. To facilitate scheduling, this program is split into three proposals. This proposal covers 352 orbits (22 weeks) from 31 August 2009 to 31 January 2010.
CCD Hot Pixel Annealing
All the data for this program is acquired using internal targets (lamps) only, so all of the exposures should be taken during Earth occultation time (but not during SAA passages). This program emulates the ACS pre-flight ground calibration and post-launch SMOV testing (program 8948), so that results from each epoch can be directly compared. Extended Pixel Edge Response (EPER) and First Pixel Response (FPR) data will be obtained over a range of signal levels for the Wide Field Channel (WFC). The High Resolution Channel (HRC) visits have been removed since it could not be repaired during SM4.
Internal FUV Wavelength Verification
This program will be executed after the uplink of the OSM1 position updates derived from the determination of the wavelength-scale zero points and desired spectral ranges for each grating in Activity COS29 (Program 11487 – COS FUV Internal/External Wavelength Scales). This program will verify that the operational spectral ranges for each grating, central wavelength, and FP-POS are those desired. Subsequent to a successful verification, COS FUV ERO observations that require accurate wavelength scales (if any) and FUV science can be enabled. An internal wavelength calibration spectrum using the default PtNe lamp (lamp 1) with each FUV grating at each central wavelength setting and each FP-POS position will be obtained for the verification. Additional exposures and waits between certain exposures will be required to avoid – and to evaluate – mechanism drifts.
COS FUV External Spectroscopic Performance – Part 1
The goal of this project is to measure the spectral resolution of absorption lines for each COS FUV grating. We will acquire science data at one central wavelength for each grating through both PSA and BOA apertures. We will evaluate the spectral resolution at offset pointings characteristic of routine COS observations. The targets chosen for these observations have sharp absorption lines. First, a target acquisition is performed to place the target at the center of the aperture. For each FUV grating we obtain moderately high S/N (~50 per resel counting statistics) observations at one central wavelength setting with the PSA. We repeat the sequence of observations at each of four additional POS-TARG pointings to form a diamond pattern whose vertices are displaced 0.25 arcsec in either X or Y from the initial pointing. The purpose of these observations is to characterize the degradation in spectral resolution due to small positioning errors in the peakup process. These observations will be obtained in Time-Tag mode and will utilize Flash=Yes. The total exposure time in all 4 offset positions is equal to the exposure time at the central position. At each offset position, there are two exposures, one each at FP positions 3 and 1. For one grating (G130M) and central wavelength (1291) only we obtain a Time-Tag exposure with Flash=No to verify auto-wavecal capability. In addition, on a brighter target, we obtain a single TTag exposure for each grating with the object centered in the BOA aperture, in order to measure the spectral resolution in this mode. These spectra will have S/N=20. Finally, for each grating we obtain one Time-Tag exposure and one ACCUM exposure with the object centered in the aperture to verify that the Doppler correction is being properly exectuted on board in ACCUM mode.
The number of exposures in this program is somewhat greater than is described in the Activity Summary, but reflects what was intended to be measured.
This is SMOV Program COS 31.
Implementation Method: Stored commanding
Dependencies: Execute after verification based upon analysis of COS 30 (Internal FUV Wavelength Verification)
Applicable SMOV Requirements: L.10.4.2.3.12.2; L.10.4.2.3.13; L.10.4.2.1.8.
CCD Dark Monitor Part 1
Monitor the darks for the STIS CCD.
CCD Bias Monitor-Part 1
Monitor the bias in the 1×1, 1×2, 2×1, and 2×2 bin settings at gain=1, and 1×1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns.
Cycle 17 STIS CCD Imaging Flats
This program periodically monitors the STIS CCD imaging mode flat fields by using the tungsten lamps.
CCD Read Noise Monitor
This proposal measures the read noise of all the amplifiers (A, B, C, D) on the STIS CCD using pairs of bias frames. Full-frame and binned observations are made in both Gain 1 and Gain 4, with binning factors of 1×1, 1×2, 2×1, and 2×2. All exposures are internals. Pairs of visits are scheduled monthly for the first four months and then bi-monthly after that.
Elucidating the Mystery of the Io Footprint Time Variations
The Io UV footprint (IFP) is an auroral emission on Jupiter consisting of one or more spots resulting from the electromagnetic interaction between Io and the Jovian magnetosphere. Recent UV HST observations of the Jovian aurora raised new issues and put previous interpretations under question. Dedicated STIS Time-tag observations based on only 3 HST orbits will help us to directly answer the following questions and test new hypothesis on the physics driving their associated phenomenon.
The proposed observations will determine whether the previously observed short timescale (~2 min) variations of the IFP are periodic or burst events. If the (quasi-) periodicity is established, these constraints will help us to understand the origin of these variations. These observations will also clarify the conditions of occurrence of the unexpected quasi-simultaneous variations of the southern multiple spots of the IFP. Moreover, we propose to observe the emergence of the southern leading (or precursor) spot and the possible evolution of its brightness. These two elements might validate or exclude the recently proposed idea that cross-hemisphere electron beams or strong non- linearities of the electromagnetic interaction explain the presence of the leading and secondary spots.
STIS-20 NUV MAMA Dark Monitor
The STIS NUV-MAMA dark current is dominated by a phosphorescent glow from the detector window. Meta-stable states in this window are populated by cosmic ray impacts, which, days later, can be thermally excited to an unstable state from which they decay, emitting a UV photon. The equilibrium population of these meta-stable states is larger at lower temperatures; so warming up the detector from its cold safing will lead to a large, but temporary, increase in the dark current.
To monitor the decay of this glow, and to determine the equilibrium dark current for Cycle 17, four 1380s NUV-MAMA ACCUM mode darks should be taken each week during the SMOV period. Once the observed dark current has reached an approximate equilibrium with the mean detector temperature, the frequency of this monitor can be reduced to one pair of darks per week.
WFC3 UVIS CCD Daily Monitor
The behavior of the WFC3 UVIS CCD will be monitored daily with a set of full-frame, four-amp bias and dark frames. A smaller set of 2Kx4K subarray biases are acquired at less frequent intervals throughout the cycle to support subarray science observations. The internals from this proposal, along with those from the anneal procedure (11909), will be used to generate the necessary superbias and superdark reference files for the calibration pipeline (CDBS).
Cycle 17: UVIS Bowtie Monitor
Ground testing revealed an intermittent hysteresis type effect in the UVIS detector (both CCDs) at the level of ~1%, lasting hours to days. Initially found via an unexpected bowtie-shaped feature in flatfield ratios, subsequent lab tests on similar e2v devices have since shown that it is also present as simply an overall offset across the entire CCD, i.e., a QE offset without any discernable pattern. These lab tests have further revealed that overexposing the detector to count levels several times full well fills the traps and effectively neutralizes the bowtie. Each visit in this proposal acquires a set of three 3×3 binned internal flatfields: the first unsaturated image will be used to detect any bowtie, the second, highly exposed image will neutralize the bowtie if it is present, and the final image will allow for verification that the bowtie is gone.
WFC3/UVIS External and Internal CTE Monitor
CCD detector Charge Transfer Inefficiency (CTI)-induced losses in photometry and astrometry will be measured using observations of the rich open cluster NGC6791 and with the EPER (Extended Pixel Edge Response) method using tungsten lamp flat field exposures. Although we do not expect to see CTE effects at the outset of Cycle 17, this CTE monitoring program is the first of a multi-cycle program to monitor and establish CTE-induced losses with time. We expect to measure CTE effects with a precision comparable to the ACS measurements.
FLIGHT OPERATIONS SUMMARY:
Significant Spacecraft Anomalies: (The following are preliminary reports of potential non-nominal performance that will be investigated.)
11996 – GSAcq(1,2,1) scheduled at 243/13:53:25z – 14:00:56 failed to fine lock backup (1,0,1) using FGS-1.
Observations possibly affected: COS 62 thru 68, Proposal ID# 11489
Subsequent REAcq(1,2,1) scheduled at 243/15:27:26z resulted in fine lock backup (1,0,1) using FGS-1.
Observations possibly affected: COS 71 thru 76 Proposal ID# 11489, ACS 26 Proposal ID# 11879.
HSTARS FOR DAY 226:
11997 – GSAcq(2,1,1) at 226/03:52:43z required two attempts to achieve FL-DV on FGS2. FGS2 initially failed a walkdown with scan step limit exceeded, but was able to achieve FL-DV on the second try. The acquisition was successful.
COMPLETED OPS REQUEST: (None)
COMPLETED OPS NOTES: (None)
FGS GSAcq 08 08
FGS REAcq 10 10
OBAD with Maneuver 07 07
SIGNIFICANT EVENTS: (None)